Which of the following is equal to the rational expression when x does not equal -4 or 16?
x^2-16/(X-16)(x+4)
A.1/x+4
B.x-4/x-16
C.x/x+7

Answer by fractalier(6550)

Factor and cancel.
x^2-16/(x-16)(x+4) = (x+4)(x-4) / (x-16)(x+4) = x-4 / x-16
Choice B.